van der Berg and Huang have reported a method for the determination of the concentration of vanadium (V) in sea water by adsorptive stripping voltammetry after forming a complex with catechol. The Vanadium-catechol complex is deposited on a hanging mercury drop electrode at a potential of -0.1 V versus a Ag/AgCl reference electrode. A cathodic potential scan gives a stripping peak that is proportional to the concentration of V. The following standard additions were used to analyze a sample of seawater. 6.
Answer #1

Since it can assumed that there is no volume change in the due to the added quantity of Vanadium, it can be said that the left column of the data table indicates the cumulative moles of added Vanadium.

For every 2 x 10-8 addition in the number of moles of Vanadium, there is change of 8.89 nA of peak current (average value calculated).

When no [V] is added, the value of peak current can be extrapolated to 15.11nA (for no addition).
Thus the peak current of 15.11 nA corresponds to the initial concentration.
By this analogy, initial concentration can be given by \frac{15.11*10^{-9} * 2*10^{-8}}{8.89*10^{-9}} = 3.4 *10 ^{-8} M